### v1

# Majority-Logic Decoding of Binary Locally Recoverable Codes: A Probabilistic Analysis

Abstract
Locally repairable codes (LRCs) were originally introduced to enable efficient recovery from erasures in distributed storage systems by accessing only a small number of other symbols. While their structural properties—such as bounds and constructions—have been extensively studied, the performance of LRCs under random erasures and errors has remained largely unexplored. In this work, we study the error- and erasure-correction performance of binary linear LRCs under majority-logic decoding (MLD). Focusing on LRCs with fixed locality and varying availability, we derive explicit upper bounds on the probability of decoding failure over the memoryless Binary Erasure Channel (BEC) and Binary Symmetric Channel (BSC). Our analysis characterizes the behavior of the bit-error rate (BER) and block-error rate (BLER) as functions of the locality and availability parameters. We show that, under mild growth conditions on the availability, the block decoding failure probability vanishes asymptotically, and that majority-logic decoding can successfully correct virtually all of error and erasure patterns of weight linear in the blocklength. The results reveal a substantial gap between worst-case guarantees and typical performance under stochastic channel models.
I Introduction
Locally repairable codes (LRCs) are a class of error-correcting codes designed to enable the recovery of individual code symbols by accessing only a small subset of other symbols, known as local recovery sets (also referred to as repair groups). This property, formalized through the notions of locality and availability, makes LRCs particularly attractive for distributed storage systems, where node failures lead to data loss modeled as erasures [10]. By providing multiple, disjoint recovery sets for each symbol, LRCs allow local repair with low bandwidth and latency overheads, and have therefore received significant attention in recent years.
Although LRCs were originally introduced to combat erasures, the existence of multiple recovery sets naturally suggests their applicability to error correction as well. In practical storage and memory systems, failures are often better modeled as random errors—such as bit flips—rather than erasures, reflecting data corruption at the level of memory elements [25]. Such bit flips arise from a variety of unavoidable physical effects and remain a major concern even in highly reliable systems, including safety-critical applications [2]. When errors are present, corrupted symbols may invalidate some recovery sets, thereby affecting the reliability of local decoding. This observation motivates the use of majority-logic decoding (MLD), a classical decoding paradigm first introduced for Reed–Muller codes [28, 20]. Under MLD, each symbol is recovered by taking a majority vote over multiple recovery sets, where each set provides an independent linear equation for the target symbol. Decoding succeeds as long as a majority (i.e., at least half) of these equations remain uncorrupted [18].
Majority-logic decoding is a simple, hard-decision decoding method with low complexity and extremely small decoding delay, which historically made it attractive for applications such as deep-space communication and memory protection [15, 16, 9]. Despite major advances in decoding algorithms and hardware capabilities, there has been a renewed interest in such low-complexity decoders for latency-critical applications, where hard-decision algorithms are often preferred over soft-decision alternatives [4]. Recently, MLD has also been connected to combinatorial designs, and to the concept of Service Rate Region, a metric that quantifies the efficiency of linear codes for optimizing data access [17]. Moreover, MLD is not a bounded-distance decoder and is known to frequently correct more errors than its guaranteed radius [6]. For certain code families, such as Grassmann codes, MLD remains the only known method enabling efficient and practical decoding [3]. Codes that admit such decoding algorithms are referred to as majority-logic decodable codes [23].
Surprisingly, despite the natural compatibility between LRCs and MLD, existing works have almost exclusively focused on the use of LRCs for erasure recovery. Even when robustness to errors is considered—for example, allowing recovery of an erased symbol in the presence of errors in parity-check symbols [24]—the explicit use of LRCs to correct random errors has not been systematically studied. In particular, no prior work has analyzed the performance of LRCs under standard probabilistic error models, such as the Binary Symmetric Channel (BSC), using MLD. Most of the existing literature on LRCs concentrates on code constructions and fundamental limits, including bounds on minimum distance, optimal length, symbol size, and channel capacity [21]. From an MLD perspective, the availability property of LRCs—namely, the presence of multiple recovery sets per symbol—naturally suggests increased robustness, as it enables majority voting across independent recovery equations and thus greater tolerance to erasures and errors. However, despite this intuition, the efficiency of LRCs against random erasures and errors—quantified in terms of decoding failure probability, bit-error rate (BER), or block-error rate (BLER)—has remained largely unexplored. In contrast, for classical code families such as Reed–Muller codes, where majority-logic decoding can be used efficiently, such probabilistic performance analyses were established quite some time ago [14, 8]. While Reed–Muller codes can be viewed as early examples of codes with local recovery structure, they are not typically designed or employed for modern distributed storage systems.
Contributions
  • MLD for error correction in LRCs: We extend the use of binary linear locally recoverable codes—originally designed for erasure recovery in distributed storage—to error correction via majority-logic decoding (MLD), a simple, practical, and low-complexity hard-decision decoder originally developed for Reed–Muller codes. In contrast to the joint-typicality decoding approach presented in [21], MLD does not require knowledge of the channel law and is well suited for latency- and complexity-constrained settings.
  • Probabilistic performance analysis under BSC and BEC: We initiate a systematic probabilistic analysis of the error- and erasure-correction performance of LRCs under MLD over the memoryless Binary Symmetric Channel (BSC) and Binary Erasure Channel (BEC). We derive explicit bounds on the probability of decoding failure as functions of the locality $r$ and availability $t$.
  • Typical-case decoding guarantees beyond worst-case bounds: Moving beyond deterministic, worst-case guarantees based on minimum distance or availability, we characterize the typical decoding behavior of LRCs under random errors and erasures. Our results quantify how many errors and erasures can be corrected with high probability using MLD and reveal a substantial gap between worst-case adversarial limits and stochastic performance.
  • Generalization of classical Reed–Muller results: Our analysis generalizes classical probabilistic decoding results for Reed–Muller codes [14] as special cases, while addressing a previously unexplored aspect of LRC performance under MLD.
This paper is organized as follows. Section II introduces binary LRCs with availability and reviews the MLD framework. Section III presents a probabilistic analysis of the decoding performance of LRCs under MLD over the BSC and BEC, including bounds on BER and BLER, and weight of correctable error and erasure patterns in the asymptotic regimes of block length $n$. Section IV provides simulation results that validate the theoretical analysis and illustrate the impact of availability scaling. Section V concludes the paper and discusses directions for future work.
In the sequel, $\log$ denotes the base-2 logarithm, $\ln$ denotes the natural logarithm, and $\exp(x)$ denotes $2^{x}$ while $\mathrm{e}$ denotes the Euler’s constant. The binary field is denoted as $\mathbb{F}_{2}$, and binary linear code $\mathcal{C}$ with parameters $n,k,d]_{2}$ is a $k$-dimensional subspace of the $n$-dimensional vector space $\mathbb{F}_{2}^{n}$ with minimum Hamming distance $d$. Set of positive integers not exceeding $i$, namely $\{1,2,\dots,i\}$, is denoted as $[i]$.
II Locally Recoverable Codes with Availability
In this section, we will introduce the Locally Recoverable Codes with Availability, and its decoding method, namely Majority-logic decoding.
II-A Locally Recoverable Codes with Availability
We follow the conventional definitions of binary linear LRCs with availability established in, for instance, [12, 29].
Definition 1.
The $i$-th code symbol of an $n,k,d]_{2}$ binary linear code $\mathscr{C}\subseteq\mathbb{F}_{2}^{n}$ is said to have locality $r$ and availability $t$ if, for every codeword $\mathbf{c}=(c_{1},c_{2},\ldots,c_{n})\in\mathscr{C}$, there exist $t$ pairwise disjoint recovery sets (also referred to as repair sets) $R^{1,i},\,R^{2,i},\,\ldots,\,R^{t,i}\subseteq[n]\setminus\{i\}$ such that:
  1. 1.
    Each set has cardinality at most $r$, i.e., $|R^{j,i}|\leq r,\ \forall\,j\in[t]$, and
  2. 2.
    For each recovery set $R^{j,i}$, the symbol $c_{i}$ can be recovered via the sum of the symbols in that set, i.e., $c_{i}=\sum\limits_{u\in R^{j,i}}c_{u},\,\forall\,j\in[t]$, where the sum is over $\mathbb{F}_{2}$.
Definition 2.
The linear code $\mathscr{C}$ is said to have all-symbol locality $r$ and availability $t$ if every code symbol $i\in[n]$ has locality $r$ and availability $t$. We refer to such a code $\mathscr{C}$ as a linear $(r,t)_{a}$-LRC.
The properties of LRCs imply that every codeword symbol can be recovered from $t$ pairwise disjoint local recovery sets, or equivalently, from $t$ independent votes. This structural recovery naturally motivates us to use majority-logic decoding (MLD), which was first introduced by Reed in 1954 for Reed–Muller codes [28, 20], where the final decision on each symbol is determined by the majority of its local votes. Codes that can be decoded using MLD are called majority-logic decodable codes. Recently, authors in [12] presented a construction of a class of long binary LRC with availability based on majority-logic decodable codes.
In this write-up, we focus on a class of binary linear LRCs $\mathscr{C}\subseteq\mathbb{F}_{2}^{n}$ in which: (1) all recovery sets have the same size $r$, where $r$ is fixed, and (2) the number of disjoint recovery sets $t$ grows with the code length $n$ (as in the case of Simplex codes). We justify that the assumption of constant recovery set size (1) is both valid and reasonable, as used commonly in the literature, for example, in [29]. Indeed, this assumption not only helps simplify our analysis but is practically solid. If a recovery set of size $s<r$ exists, it can always be formally treated as a set of size $r$ by add into it other rsr-s available symbols among the remaining. Since these extra objects do not participate in the actual recovery process, the (pairwise) disjointness of the recovery sets is preserved. Moreover, in hardware/VLSI (Very-Large-Scale Integration) systems, it is significantly more efficient to design, implement, and maintain circuits with a uniform number of inputs rather than supporting varying input sizes [11], especially as the number of required circuits scales. Such codes are often referred to as LRCs with large availability and have attracted significant practical interest [13]. Explicit constructions of binary linear LRCs with fixed locality and availability growing with $n$ have been presented, for example, in [13] (linear growth) and [22, 29]. Authors in [27] presented a construction of a class of LRC with scaling locality and availability that required large field sizes. Authors in [30] presented a construction of a class of binary code that achieves arbitrary locality and availability. Interested readers are referred to [26] for a comprehensive and detailed review.
II-B Majority-Logic Decoding
Suppose the original codeword $\mathbf{c}$ is transmitted over a communication channel and the received (possibly corrupted) sequence is denoted by $\mathbf{u}=\mathbf{c}+\mathbf{e}$, where $\mathbf{u}=(u_{1},u_{2},\ldots,u_{n})\in\mathbb{F}_{2}^{n}$, and $\mathbf{e}\in\mathbb{F}_{2}^{n}$ is a binary noise vector. For a fixed $i$ and let $R^{j,i}$ be one of the recovery sets for symbol $c_{i}$, the decoder forms a local estimate $\hat{c}_{i}^{(j)}$ of $c_{i}$ using
$\hat{c}_{i}^{(j)}=\sum_{h\in R^{j,i}}u_{h},\quad\forall\,j\in[t].$
In practice, these summations over $\mathbb{F}_{2}$ can be implemented efficiently using XOR gates. Finally, a global decision on $c_{i}$ is made by taking the majority vote among the local estimates:
[…]
Example 1.
Consider a $7,\,3]_{2}$ Simplex code whose length-$7$ codewords $\mathbf{c}=(c_{1},c_{2},c_{3},c_{4},c_{5},c_{6},c_{7})$ lie in the row space of the generator matrix
$\mathbf{G}=\begin{bmatrix}0&0&0&1&1&1&1\\ 0&1&1&0&0&1&1\\ 1&0&1&0&1&0&1\end{bmatrix}$.
In this code, the symbol $c_{1}$ admits three disjoint recovery sets and can be written as
$c_{1}=c_{2}+c_{3}=c_{4}+c_{5}=c_{6}+c_{7}\quad(\text{over }\mathbb{F}_{2}).$ (1)
Thus, $\{c_{2},c_{3}\}$, $\{c_{4},c_{5}\}$, and $\{c_{6},c_{7}\}$ form three independent recovery sets for $c_{1}$.
Suppose the received vector $\mathbf{u}$ differs from the transmitted codeword $\mathbf{c}$ due to channel errors. Since the recovery sets in (1) are pairwise disjoint, each bit error affects at most one recovery set. For example, if only $u_{2}$ is flipped while all other symbols are received correctly, then the estimate $u_{2}+u_{3}$ disagrees with $u_{4}+u_{5}$ and $u_{6}+u_{7}$. A majority vote among
$u_{2}+u_{3},\quad u_{4}+u_{5},\quad u_{6}+u_{7}$
correctly recovers $c_{1}$. Moreover, even if two bit errors occur within the same recovery set—for instance, $u_{2}=c_{2}+1$ and $u_{3}=c_{3}+1$—the corresponding estimate remains correct:
$u_{2}+u_{3}=(c_{2}+1)+(c_{3}+1)=c_{2}+c_{3}\quad(\text{over }\mathbb{F}_{2}).$
In contrast, an error pattern of the same weight affecting two distinct recovery sets (e.g., errors at $u_{2}$ and $u_{4}$) leads to an incorrect majority vote. This illustrates that a recovery set produces an incorrect vote if and only if it contains an odd number of errors.
In general, a local recovery set containing an even number of errors yields a correct vote, whereas a recovery set containing an odd number of errors yields an incorrect vote. Thus, the correctness of each vote depends only on the parity of the number of errors within the recovery set, not on the total number of errors. Consequently, majority-logic decoding can successfully correct certain error patterns that conventional bounded-distance decoding methods cannot, despite relying solely on local parity constraints. This observation motivates our probabilistic analysis of majority-logic decoding for $(r,t)_{a}$ LRCs over channels with independent bit flips and erasures, in terms of BER, BLER, and the number of correctable errors and erasures.
III Analysis of Decoding Capability of LRCs
In this section, we present two complementary approaches for analyzing the error- and erasure-correction capability of LRCs under majority-logic decoding (MLD): a deterministic approach and a probabilistic approach. The former characterizes worst-case (adversarial) correction guarantees, while the latter focuses on typical performance when errors and erasures occur randomly and independently. Our main results are derived under this probabilistic channel model.
III-A Deterministic versus Probabilistic Perspectives
Deterministic approach.
Typically, the correction capability of an LRC is characterized via minimum distance or availability under an adversarial (i.e., worst case) error model. In the worst case, even if t1t-1 recovery sets are completely lost (for example, each containing at least one erasure), the symbol can still be recovered using the remaining intact set. Hence, deterministically, an $(r,t)_{a}$-LRC can tolerate up to t1t-1 erasures, irrespective of the channel statistics [31]. This represents the worst-case or adversarial erasure-correction capability of the code. Similarly, the code can tolerate at most $\lfloor(t-1)/2\rfloor$ errors, since in this case the majority of local parity votes remain correct, allowing MLD to successfully recover the target symbol.
However, such worst-case bounds are often pessimistic and not particularly informative in practice. In contrast, a more common and realistic scenario is one in which errors and erasures occur randomly [21]. Under such stochastic behavior, errors often affect multiple symbols within the same recovery set, leaving other recovery sets completely intact. This sharply contrasts with the adversarial scenario, where errors are distributed to corrupt as many recovery sets as possible. Consequently, the number of corrupted recovery sets is often significantly smaller than the total number of raw errors, allowing successful decoding even beyond worst-case guarantees. This explains why MLD frequently corrects more than $\lfloor(t-1)/2\rfloor$ errors in practice and motivates the probabilistic analysis that follows.
Probabilistic approach.
A more realistic view assumes that erasures or bit-flips occur independently according to a stochastic channel law, such as the memoryless $\mathrm{BEC}$ or $\mathrm{BSC}$. Under this model, we analyze the probability of decoding failure at both the symbol (BER) and block (BLER) levels as functions of the channel parameters (bit-flip probability $p_{f}$ or erasure probability $p_{e}$) and the code parameters $(n,r,t)$.
III-B Binary Symmetric Channel ($\mathrm{BSC}$)
In this case, each bit is independently flipped with probability $p_{f}$ (“data corruption”) where $0\leq p_{f}<0.5$. We now characterize the symbol/bit decoding failure probability of MLD in this setting.
Theorem 1.
Consider transmission over a $\mathrm{BSC}(p_{f})$ with $\mathrm{i.i.d.}$ errors, where $0\leq p_{f}<0.5$ is a fixed bit-flip probability. For binary linear $(r,t)_{a}$-LRC, the symbol (or bit) decoding failure probability of MLD satisfies
$P^{\mathrm{BSC}}_{\mathrm{fail,bit}}$ $\leq\bigl(2\sqrt{a(1-a)}\bigr)^{t}=\big(1-(1-2p_{f})^{2r}\big)^{t/2},$
$\quad\text{where}\quad a=\frac{1+(1-2p_{f})^{r}}{2}.$
Proof.
For a given recovery set $R^{x}$ of size $r$, let
$W$ $\triangleq\Pr[R^{x}\text{ gives a wrong vote}]$
$=\Pr[R^{x}\text{ contains an odd number of bit flips}]$
$=\sum_{\text{odd }s}\binom{r}{s}p_{f}^{s}(1-p_{f})^{r-s},$
and
$C$ $\triangleq\Pr[R^{x}\text{ gives a correct vote}]$
$=\Pr[R^{x}\text{ contains an even number of bit flips}]$
$=\sum_{\text{even }s}\binom{r}{s}p_{f}^{s}(1-p_{f})^{r-s}.$
Clearly, $W+C=1$, and by the Binomial theorem expansion,
$C-W$ $=\sum_{\text{even }s}\binom{r}{s}p_{f}^{s}(1-p_{f})^{r-s}-\sum_{\text{odd }s}\binom{r}{s}p_{f}^{s}(1-p_{f})^{r-s}$
$=((1-p_{f})-p_{f})^{r}=(1-2p_{f})^{r}.$
Hence,
$C=\frac{1+(1-2p_{f})^{r}}{2},\qquad W=\frac{1-(1-2p_{f})^{r}}{2}.$ (2)
Note that since $0\leq p_{f}<0.5$, we have $W<\frac{1}{2}<C$.
Random-variable formulation. For each recovery set $R^{x}$, define
[…]
Then $\Pr[\xi_{x}=1]=C$ and $\Pr[\xi_{x}=-1]=W$, consequently
$\mathbb{E}[\xi_{x}]=C-W=(1-2p_{f})^{r}=y^{r},\qquad\mathrm{Var}(\xi_{x})=1-y^{2r},$
where $y\triangleq 1-2p_{f}$. MLD fails when the total number of wrong votes exceeds the correct ones, i.e.,
$\sum_{x=1}^{t}\xi_{x}\leq 0.$
Note that different $\xi_{x}$ are obtained from pairwise disjoint recovery sets $R^{x}$ and therefore independent. Thus, the decoding error probability is $P^{\mathrm{BSC}}_{\mathrm{fail,bit}}=\Pr\!\left(\sum_{x=1}^{t}\xi_{x}\leq 0\right).$ The result follows immediately using Lemma 1. ∎
Lemma 1 (Chernoff bound).
Let $X_{1},\dots,X_{t}$ be $\mathrm{i.i.d.}$ random variables taking values in $\{-1,1\}$, with $\Pr(X_{i}=1)=a$ and $\Pr(X_{i}=-1)=1-a$. If $a>1/2$, then for any $t\geq 1$, the cumulative sum $S_{t}=\sum_{i=1}^{t}X_{i}$ satisfies
$\Pr(S_{t}\leq 0)\leq{\left(2\sqrt{a(1-a)}\right)^{t}}.$
Proof.
The result follows from the standard exponential moment method (see, e.g., [5, Ch. 2]). For any $\lambda<0$, Markov’s inequality applied to the random variable $e^{\lambda S_{t}}$ implies
$\Pr(S_{t}\leq 0)=\Pr(e^{\lambda S_{t}}\geq 1)\leq\mathbb{E}[e^{\lambda S_{t}}].$
By independence, the moment generating function factorizes as $\mathbb{E}[e^{\lambda S_{t}}]=\prod_{i=1}^{t}\mathbb{E}[e^{\lambda X_{i}}]=(ae^{\lambda}+(1-a)e^{-\lambda})^{t}$. We minimize the base term $\phi(\lambda)=ae^{\lambda}+(1-a)e^{-\lambda}$ with respect to $\lambda$. Setting the derivative to zero yields the optimal choice $e^{\lambda}=\sqrt{(1-a)/a}$. Since $a>1/2$, we have $e^{\lambda}<1$, ensuring $\lambda<0$ as required. Substituting this value back into $\phi(\lambda)$ yields the bound $\big(2\sqrt{a(1-a)}\big)^{t}$. ∎
Remark 1 (Exponential Tightness).
The bound is exponentially tight; that is, the base of the exponential term cannot be improved. This follows from Cramér’s theorem in the theory of Large Deviations (see [7, Ch. 2]). The rate of decay is governed by the Kullback–Leibler divergence between a Bernoulli(1/2)(1/2) and a Bernoulli(a)(a) distribution:
$\lim_{t\to\infty}-\frac{1}{t}\ln\Pr(S_{t}\leq 0)=D_{\mathrm{KL}}(1/2\,\|\,a)$
$=\frac{1}{2}\ln\left(\frac{1/2}{a}\right)+\left(1-\frac{1}{2}\right)\ln\left(\frac{1-1/2}{1-a}\right)$
$=-\ln\left(2\sqrt{a(1-a)}\right).$
Consequently, $\Pr(S_{t}\leq 0)\approx\mathrm{e}^{-t\cdot D_{\mathrm{KL}}(1/2\,\|\,a)}=\left(2\sqrt{a(1-a)}\right)^{t}$. This confirms that both the base and the linear dependence on $t$ are asymptotically optimal. Note that although the base of the exponent is optimal, the bound can be further refined by incorporating the polynomial pre-factor established by the Bahadur-Rao theorem [1].
Remark 2.
Theorem 1 does not assume how $t$ scales with $n$. In particular, the bound holds uniformly for all values of $n$ and $t$, regardless of whether $t$ is fixed, grows slowly, or grows unbounded with $n$.
On the other hand, it follows from (2) that the quantity CC—the probability that a given recovery set produces a correct vote—is a decreasing function of the locality $r$ for a fixed bit-flip probability $p_{f}\in[0,\,0.5)$. This behavior may appear counterintuitive, since $C$ depends only on the parity (even or odd) of the number of bit flips within the recovery set. This implies that, when the availability $t$ is fixed, smaller recovery sets yield a higher probability of a correct vote.
Theorem 1 leads to the following corollary on the block decoding failure probability (BLER).
Theorem 2 (MLD Success for LRCs on the BSC).
Consider a binary $(r,t)_{a}$-LRC where $r$ is fixed and $t=t(n)$ grows with $n$. When transmitted over a $\mathrm{BSC}(p_{f})$ with fixed error probability $0\leq p_{f}<0.5$, the code achieves successful block decoding asymptotically if the availability scales sufficiently fast. Concretely, if $t(n)=\omega(\log n)$ (i.e., $t(n)$ grows faster with $n$ than $\log n$ or $\lim\limits_{n\to\infty}\frac{t(n)}{\log n}=\infty$), then the probability of block decoding failure vanishes in the asymptotic regime of $n$:
$\lim\limits_{n\rightarrow\infty}P_{\mathrm{fail,block}}^{\mathrm{BSC}}=0.$
Proof.
For any fixed symbol, Theorem 1 gives the bit-failure probability
$P_{\mathrm{fail,bit}}^{\mathrm{BSC}}\leq\bigl(1-(1-2p_{f})^{2r}\bigr)^{t/2}.$
Using the union bound, the failure probability of decoding blocks (or words) satisfies
$P_{\mathrm{fail,block}}^{\mathrm{BSC}}\;\leq\;n\,P_{\mathrm{fail,bit}}^{\mathrm{BSC}}\leq n\bigl(1-(1-2p_{f})^{2r}\bigr)^{t/2}.$
Define
$x\triangleq(1-2p_{f})^{2r}\in(0,1],$
so that the upper bound becomes $n(1-x)^{t/2}$. Using the standard inequality $\log(1-x)<-x$ for all $x\in(0,1)$, we may write
$\log(1-x)=-x-\varepsilon,\qquad\text{for some }\varepsilon>0.$
Thus
$n(1-x)^{t/2}=\exp\!\left(\log n-\frac{t}{2}x-\frac{t}{2}\varepsilon\right).$
Since $x=(1-2p_{f})^{2r}$ and $\varepsilon$ are both positive, the exponent tends to -$\infty$ provided
$\lim_{n\to\infty}\frac{t(n)}{\log n}\;\longrightarrow\;\infty.$
In this case we have $n(1-x)^{t/2}\to 0$, and therefore $P_{\mathrm{fail,block}}^{\mathrm{BSC}}\to 0$. This completes the proof. ∎
The result above leads to the following interesting result.
Theorem 3 (MLD of LRCs with availability $t$ and locality $r$ over the $\mathrm{BSC}$).
Let $c\geq 3$ be any constant. For a binary $(r,t)_{a}$-LRC with availability $t=t(n)$ growing with $n$, all but a vanishing fraction of error patterns of weight
[…]
are correctly recovered by majority-logic decoding as $n\to\infty$. In other words, virtually all error patterns of weight $w$ are correctable in the asymptotic regime of $n$.
Proof.
Consider a $\mathrm{BSC}(p)$ channel in which
$p\triangleq\frac{1}{2}\!\left(1-\sqrt[2r]{\frac{c\log n}{t}}\right),\qquad x\triangleq(1-2p)^{2r}=\frac{c\log n}{t},$
so that error patterns of weight $w=pn$ correspond to the threshold in the right-hand side of (3). Using Union bound and Theorem 1, the block decoding failure probability under majority-logic decoding satisfies
$P_{\mathrm{fail,block}}^{\mathrm{BSC}}\leq nP_{\mathrm{fail,bit}}^{\mathrm{BSC}}\;\leq\;n(1-x)^{t/2}.$
Using $\log(1-x)=-x-\varepsilon$ for some $\varepsilon>0$ and $x\in(0,1)$,
$P_{\mathrm{fail,block}}^{\mathrm{BSC}}$ $\leq\exp\!\left(\log n-\frac{t}{2}x-\frac{t}{2}\varepsilon\right)$
$=\exp\!\left(\left(1-\frac{c}{2}\right)\log n-\frac{t}{2}\varepsilon\right)\ (\text{since }tx=c\log n.)$
Now consider the set of weight-$pn$ erasure patterns,
$A_{m}=\{\mathbf{e}\in\{0,1\}^{n}:\mathrm{wt}(\mathbf{e})=pn\}.$
By the classical lower bound on binomial probabilities [19, Ch. 10, Lem. 7],
$P(A_{m})$ $=\binom{n}{pn}p^{pn}(1-p)^{(1-p)n}\geq\frac{1}{\sqrt{8np(1-p)}}$
$\geq\frac{1}{\sqrt{2n}}\hskip 18.06749pt\text{($4p(1-p)\leq 1$ by AM-GM inequality)}$
$=\Theta(n^{-1/2})=\exp\!\left(-\frac{1}{2}\log n+o(1)\right).$
Let $B_{m}\subseteq A_{m}$ be the subset of weight-$pn$ patterns on which MLD fails. Because all weight-$pn$ erasure patterns have identical probability under the $\mathrm{BSC}$ (i.e., they have the same type [Csiszár_Körner_2011, Ch. 2]),
$P(A_{m})$ $=|A_{m}|p^{pn}(1-p)^{(1-p)n},$
$P(B_{m})$ $=|B_{m}|p^{pn}(1-p)^{(1-p)n}.$
Thus the fraction of uncorrectable weight-$pn$ patterns is
$\frac{|B_{m}|}{|A_{m}|}=\frac{P(B_{m})}{P(A_{m})}.$
Since any block failure must occur on some error pattern,
$P(B_{m})\leq P_{\mathrm{fail,block}},$
and therefore
$\frac{|B_{m}|}{|A_{m}|}=\frac{P(B_{m})}{P(A_{m})}\leq\frac{P_{\mathrm{fail,block}}}{P(A_{m})}.$
Substituting the upper bound for $P_{\mathrm{fail,block}}^{\mathrm{BSC}}$ and the lower bound for $P(A_{m})$,
$\frac{|B_{m}|}{|A_{m}|}\leq\exp\!\left(\frac{3}{2}\log n-\frac{c}{2}\log n-\frac{t}{2}\varepsilon-o(1)\right).$
If $c\geq 3$ and $t(n)\to\infty$, the exponent tends to -$\infty$, and hence
$\frac{|B_{m}|}{|A_{m}|}\longrightarrow 0.$
Thus, asymptotically, only a vanishing fraction of weight-$pn$ error patterns are uncorrectable, proving that virtually all such patterns are correctly decoded. ∎
Remark 3.
The constant $c$ in  (4) and (III-C) captures a trade-off between the weight of the correctable error patterns and the rate at which the fraction of uncorrectable patterns decays. Choosing a smaller value of $c$ increases the threshold weight $w$ in (4), allowing MLD to correct error patterns of higher weight. However, this comes at the cost of a slower decay in the upper bound (III-C) on the fraction of uncorrectable patterns. Conversely, larger values of $c$ yield faster decay of this fraction but guarantee correctability only for error patterns of smaller weight.
On the other hand, having larger values of $t$ helps to yield faster decay on the fraction of uncorrectable patterns, and also guarantee correctability for error patterns of higher weight.
Remark 4.
Unlike Theorem 2, Theorem 3 does not impose any condition on the growth rate of $t(n)$. However, if $t(n)=\omega(\log n)$ (as in the setting of Theorem 2), then the weight threshold
[…]
satisfies $\lim_{n\to\infty}\frac{w}{n}=\frac{1}{2}$. In this regime, MLD can asymptotically correct virtually all error patterns of weight up to half the entire blocklength $n$. In particular, the typical error patterns corrected by MLD may have weight far exceeding $\frac{d_{\min}}{2}$ (i.e., half the minimum distance of the underlying code)—a phenomenon that reflects the large gap between worst-case (deterministic) and typical-case (probabilistic) error correction capability.
In contrast, Theorem 2 does not explicitly define a weight threshold for correctable error patterns. Instead, relying on the assumption of a bit flip probability $p_{f}\in[0,0.5)$, the Law of Large Numbers implies that the weight of the error patterns in a $\mathrm{BSC}(p_{f})$ channel will concentrate around $np_{f}<n/2$, which is consistent with the asymptotic bound derived in Theorem 3.
Remark 5 (Numerical example).
For $n=2^{10}$, $t(n)=\log^{2}n$, $r=4$, and $c=2$:
$\log n=\log 1024$ $=10,\quad t=100,$
$\left(\frac{c\log n}{t}\right)^{\!1/8}$ $=\left(\frac{2}{10}\right)^{\!1/8}\approx 0.818.$
Thus
$w\approx 2^{10}\,(1-0.818)\approx 186.$
So with $t=\log^{2}n$, MLD succeeds for virtually all error patterns up to $w\approx 186$. By contrast, the worst-case deterministic availability guarantee only ensures recovery from up to $\lfloor\frac{t-1}{2}\rfloor=49$ errors (e.g., if each hits a different recovery set). There is a huge difference in decoding performance between the typical-case and the worst-case scenarios.
III-C Binary Erasure Channel ($\mathrm{BEC}$)
In this case, the communication channel is a $\mathrm{BEC}$ where each bit is independently erased with probability $p_{e}$ (“data loss”) where $0\leq p_{e}<1$.
For a given recovery set $R^{x,k}$ of size $r$:
$\Pr[\text{no erasure in }R^{x,k}]=(1-p_{e})^{r}.$
Since the $t$ recovery sets are pairwise disjoint (hence independent), the probability that all sets contain at least one erasure is
$\Pr[\text{all }R^{x,k}\text{ have $\geq 1$ erasure}]=\big(1-(1-p_{e})^{r}\big)^{t}.$
This leads us to the following result.
Proposition 1.
The probability of symbol/bit decoding failure (all sets contains at least one erased symbol) over a $\mathrm{BEC}$ where bits are erased independently with probability $p_{e}$ is
$P_{\mathrm{fail,bit}}^{\mathrm{BEC}}=\big(1-(1-p_{e})^{r}\big)^{t}.$
Since $0\leq 1-(1-p_{e})^{r}<1$, the probability of decoding failure decays as $t$ grows. This observation leads to the following interesting result.
Theorem 4 (MLD for LRCs with availability $t$ and locality $r$ over the $\mathrm{BEC}$).
Consider a binary $(r,t)_{a}$-LRC where $r$ is fixed and $t=t(n)$ grows with $n$. Suppose the codeword is transmitted over a $\mathrm{BEC}(p_{e})$ with $\mathrm{i.i.d.}$ erasures, where $0\leq p_{e}<1$ is fixed. If $t(n)$ grows faster than $\log n$ as $n$ tends to infinity, then the probability of block decoding failure vanishes in the asymptotic regime of $n$:
$\lim\limits_{n\rightarrow\infty}P_{\mathrm{fail,block}}^{\mathrm{BEC}}=0.$
Proof.
The proof is similar to that of Theorem 2. We present it here for completeness.
For any fixed symbol, Proposition 1 gives the erasure decoding failure probability
$P_{\mathrm{fail,bit}}^{\mathrm{BEC}}=\big(1-(1-p_{e})^{r}\big)^{t}.$
Using the union bound, the failure probability of decoding blocks (or words) satisfies
$P_{\mathrm{fail,block}}^{\mathrm{BEC}}\;\leq\;n\,P_{\mathrm{fail,bit}}^{\mathrm{BEC}}=n\big(1-(1-p_{e})^{r}\big)^{t}.$
Define
$x\triangleq(1-p_{e})^{r}\in(0,1],$
so that the upper bound becomes $n(1-x)^{t}$. Using the standard inequality $\log(1-x)<-x$ for all $x\in(0,1)$, we may write
$\log(1-x)=-x-\varepsilon,\qquad\text{for some }\varepsilon>0.$
Thus
$n(1-x)^{t}=\exp\!\left(\log n-tx-t\varepsilon\right).$
Since $x=(1-p_{e})^{r}$ and $\varepsilon$ are both positive, the exponent tends to -$\infty$ provided
$\lim_{n\to\infty}\frac{t(n)}{\log n}\;\longrightarrow\;\infty.$
In this case we have $n(1-x)^{t}\to 0$, and therefore $P_{\mathrm{fail,block}}^{\mathrm{BEC}}\to 0$. This completes the proof. ∎
Following the same spirit of Theorem 3, one obtains the following result.
Theorem 5.
Let $c\geq\frac{3}{2}$ be any constant. For a binary $(r,t)_{a}$-LRC with availability $t=t(n)$ growing with $n$, all but a vanishing fraction of erasure patterns of weight
$w\;\leq\;n\!\left(1-\sqrt[r]{\frac{c\log n}{t}}\right)$ (4)
are correctly recovered by majority-logic decoding as $n\to\infty$. In other words, virtually all erasure patterns of weight $w$ are correctable in the asymptotic regime of $n$.
Proof.
The proof follows the same structure as that of Theorem 3; we therefore highlight only the essential differences. Set
$p$ $\triangleq 1-\sqrt[r]{\frac{c\log n}{t}},\qquad\text{so }\ 0\leq p\leq 1\$
$\text{ and }\ x$ $\triangleq(1-p)^{r}=\frac{c\log n}{t}.$
Using Union bound and Proposition 1, one has the block-failure bound
$P_{\mathrm{fail,block}}^{\mathrm{BEC}}\;\leq\;n(1-x)^{t}.$
Using the identity $\log(1-x)=-x-\varepsilon$ for some $\varepsilon>0$ when $x\in(0,1)$,
$n(1-x)^{t}$ $=\exp\!\bigl(\log n-tx-t\varepsilon\bigr)$
$=\exp\!\bigl((1-c)\log n-t\varepsilon\bigr)\qquad(\text{since }tx=c\log n.)$
Thus
$P_{\mathrm{fail,block}}^{\mathrm{BEC}}\leq\exp\!\bigl((1-c)\log n-t\varepsilon\bigr).$
Let $A_{m}$ denote the set of all error patterns of weight $pn$, and let $B_{m}\subseteq A_{m}$ be the subset on which MLD fails. As in the proof of Theorem 3, all patterns in $A_{m}$ have identical probability under the $\mathrm{BSC}$, and therefore
$\frac{|B_{m}|}{|A_{m}|}=\frac{P(B_{m})}{P(A_{m})}\leq\frac{P_{\mathrm{fail,block}}}{P(A_{m})}.$
Moreover, by the classical binomial lower bound [19, Ch. 10, Lem. 7],
$P(A_{m})\geq\Theta(n^{-1/2})=\exp\!\left(-\frac{1}{2}\log n+o(1)\right).$
Substituting the upper bound for $P_{\mathrm{fail,block}}^{\mathrm{BEC}}$ and the lower bound for $P(A_{m})$,
$\frac{|B_{m}|}{|A_{m}|}$ $\;\leq\;\frac{\exp\!\bigl((1-c)\log n-t\varepsilon\bigr)}{\exp\!\left(-\frac{1}{2}\log n+o(1)\right)}$
$=\exp\!\left(\frac{3}{2}\log n-c\log n-t\varepsilon-o(1)\right).$ (5)
Because $c\geq\frac{3}{2}$ and $\varepsilon>0$, the exponent tends to -$\infty$ when $t(n)$ tends to $\infty$ with $n$, and thus
$\frac{|B_{m}|}{|A_{m}|}\;\longrightarrow\;0.$
Hence asymptotically, only a vanishing fraction of weight-$pn$ erasure patterns are uncorrectable, proving that virtually all such patterns are correctly decoded. ∎
Remark 6.
Comparing Theorems 4 and 1, we observe that the BER over the $\mathrm{BEC}$ decays at twice the exponential rate of the corresponding failure probability over the $\mathrm{BSC}$. Indeed, BER of the $\mathrm{BEC}$ behaves like $(1-k_{1})^{t}$, whereas that of the $\mathrm{BSC}$ behaves like $(1-k_{2})^{t/2}$, where
$k_{1}=(1-p_{e})^{r}\qquad\text{and}\qquad k_{2}=(1-2p_{f})^{2r}$
are fixed constants. Thus, majority-logic decoding is exponentially more robust to erasures than to bit flips. This two-fold difference in decay rates manifests itself in the asymptotic gap between the number of correctable erasures in Theorem 5 (on the order of $n$) and the number of correctable errors in Theorem 3 (on the order of $n/2$). This observation is consistent with the classical principle in coding theory that codes can typically correct twice as many erasures as errors.
IV Simulations
To validate the theoretical bounds derived for MLD over the BSC, we performed simulations analyzing both BER and BLER. The simulations utilized a fixed locality parameter of $r=4$ while varying the availability $t(n)$ across different growth regimes relative to the block length $n$. The plotted numerical values are each averaged over $5\times 10^{6}$ runs.
IV-A Bit Decoding Failure Probability
Figure 1 compares the empirical bit decoding failure probability (BER) against the theoretical upper bounds established in Theorem 1 for a BSC with bit-flip probability $p_{f}=0.2$. We examined three distinct regimes for the availability $t$: linear growth ($t=n/4$), polylogarithmic growth ($t=(\log_{2}n)^{2}$), and sub-logarithmic growth ($t=\sqrt{\log_{2}n}$).
The results demonstrate a clear threshold behavior dependent on the growth rate of $t$:
  • Linear and Polylogarithmic Regimes: In cases where $t$ grows linearly or as $(\log_{2}n)^{2}$, the empirical failure probability decays rapidly to zero as $n$ increases, confirming that MLD successfully corrects errors when the number of recovery sets is sufficiently large. Notably, the empirical performance is consistently superior to the theoretical upper bounds, though the gap narrows as $n$ grows. The tightness of the bound is directly related to the growth rate of $t(n)$; for the linear case ($t=n/4$), the theoretical bound closely approximates the empirical probability for $n\geq 2^{12}$, whereas it provides a looser estimate in the slower-growth regimes (e.g., $t=\sqrt{\log n}$). Note that as mentioned in Remark 1, although being exponentially tight, the theoretical upper bound can be further refined by incorporating a multiplicative pre-factor.
  • Sub-logarithmic Regime: Conversely, for the slowly growing availability $t=\sqrt{\log_{2}n}$, the decoding failure probability does not vanish. Instead, it saturates at a high error floor (approximately 0.2 to 0.4), indicating that when availability grows slower than logarithmic factors, the majority vote may fail to correct errors reliably.
IV-B Block Decoding Failure Probability
Figure 2 focuses on the upper bound of the BLER, derived via the union bound on the Bit Error Rate (BER) (i.e., $P_{\mathrm{fail,block}}\leq nP_{\mathrm{fail,bit}}$) for a BSC with $p_{f}=0.13$. Since this BLER is derived from a union bound, the resulting values may exceed both the true BLER and even $1$, which is reflected in the vertical axis of the plot. We do not report empirical BLER values, as their computation becomes infeasible for large blocklengths: verifying correct block decoding requires checking the correctness of all symbols, each of which involves an increasing number of recovery sets as $n$ grows. This figure provides a fine-grained analysis of the polylogarithmic regime by varying the exponent of the logarithmic growth, specifically $t=(\log_{2}n)^{\alpha}$ for $\alpha\in\{1.8,1.9,2.05\}$. Figure 2 highlights the sensitivity of block-level decoding success to the exponent $\alpha$:
  • Failure Regime ($\alpha=1.8$): For $t=(\log_{2}n)^{1.8}$, the bound on BLER initially rises and remains high, failing to converge to 0 until $n\geq 2^{23}$. This suggests that this growth rate is insufficient to ensure vanishing BLER at practical block lengths.
  • Success Regimes ($\alpha=1.9,2.05$): As the exponent increases, the failure probability transitions into a clear decay phase, vanishing significantly faster than in the previous case. This confirms that for block-level reliability—which requires every single bit to be correct—the availability must scale with a sufficiently high power of $\log n$ to counteract the cumulative bit error probability resulting from the increasing block length.
The simulations confirm that while MLD is robust, its asymptotic reliability on the BSC is strictly governed by the scaling law of the availability $t$. Linear availability provides the strongest protection, whereas polylogarithmic availability requires a sufficient growth exponent to ensure successful block decoding.
Figure 1: Bit Decoding Failure Probability: Empirical vs Theoretical ($p_{f}=0.2,r=4$). The plot compares linear, polylogarithmic, and sub-logarithmic availability regimes.
Figure 2: Upper (Union) bound for Block Decoding Failure ($p_{f}=0.13,r=4$). The plot demonstrates the impact of the logarithmic exponent $\alpha$ on block decoding success.
V Conclusion
We analyzed the performance of majority-logic decoding for binary linear locally recoverable codes over stochastic channels. Our results show that, with sufficiently large availability, LRCs achieve vanishing block error probability and can correct a linear fraction of random errors and erasures, despite much weaker worst-case guarantees. This highlights the effectiveness of majority-logic decoding in exploiting the availability structure of LRCs and motivates further study of probabilistic decoding behavior for other locally recoverable code families, especially for $q$-ary codes.
Acknowledgment
This work was supported in part by the NSF-BSF grant FET212026. The authors thank Alexander Barg for providing them with an original manuscript in Russian and its English translation.